OpenStudy (anonymous):

Solve 2cos^2x + cosx − 1 = 0 for x over the interval [0, 2 pi ).

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Lol, I have no idea.

OpenStudy (anonymous):

ok XD

OpenStudy (anonymous):

Let \[x = \cos \theta\] So you get \[2x^2+x-1=0\] can you finish it off ;)

OpenStudy (anonymous):

But note, you don't want x, you want theta, so you have to remember your ratios and solve for cos theta = ... what ever you get.

OpenStudy (anonymous):

is it -1,1/2

OpenStudy (anonymous):

Yes, now you can solve for theta between your interval.

OpenStudy (anonymous):

that makes no sense to me

OpenStudy (anonymous):

Mhm? Now let cos theta = -1, and 1/2

OpenStudy (anonymous):

since we have cos theta = x

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

Can you be more specific

OpenStudy (anonymous):

i dont know what you are talking about at all, all this cos jazz is confusing me

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Confusion may come from your equation saying x, and me using theta, just switch it around, shouldn't matter

OpenStudy (anonymous):

i dont get it ....

OpenStudy (anonymous):

...

OpenStudy (anonymous):



OpenStudy (anonymous):